Status Report

NASA ISS On-Orbit Status 16 May 2009

By SpaceRef Editor
May 16, 2009
Filed under , , ,
NASA ISS On-Orbit Status 16 May 2009
NASA ISS On-Orbit Status 16 May 2009

All ISS systems continue to function nominally, except those noted previously or below. Saturday — light-duty day for the ISS crew.

The crew performed the regular weekly three-hour task of thorough station cleaning. [“Uborka”, usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, Padalka conducted regular maintenance inspection & cleaning on fan screens in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9), plus dust filter replacement in the FGB.

Gennady also performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Additionally, the CDR did the periodic maintenance of the Russian POTOK-150MK (150 micron) air purification subsystems in the SM & FGB by cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting the POTOK in automatic mode.

Mike performed the daily procedure of flushing the PWD (Potable Water Dispenser) ambient line with ~50mL of water (into a towel/Ziploc bag). PWD water is currently cleared only for hygienic use, pending Russian approval.

The FE-1 also conducted the daily status check on the BCAT-4 (Binary Colloidal Alloy Test 4) science payload, running by itself since 5/12 on Sample 5. [The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour.]

For today’s VolSci (Voluntary Science) program, Wakata & Barratt reviewed procedures, then activated the MSG (Microgravity Science Glovebox) and conducted another session with the SPICE (Smoke Point In Co-flow Experiment) payload, performed in the MSG and controlled by its A31p laptop with SPICE microdrives. The activities were photographed and video-recorded. [SPICE determines the point at which gas-jet flames (similar to a Butane-lighter flame) begin to emit soot (dark carbonaceous particulate formed inside the flame) in microgravity. Studying a soot-emitting flame is important in understanding the ability of fires to spread and in control of soot in practical combustion systems space. For each test point, the crewmembers adjusted the fuel flow rate until the flame was at the smoke point. Science data are collected by still images and downlinked video.]

Gennady performed the frequent status check on the Russian BIO-5 Rasteniya-2 (“Plants-2”) experiment, verifying proper operation of the BU Control Unit and MIS-LADA Module fans (testing their air flow by hand). [Rasteniya-2 researches growth and development of plants under spaceflight conditions in the LADA-15 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP).]

At ~9:10am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

At ~4:10pm, Mike Barratt has his weekly PFC (Private Family Conference) scheduled, via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR) and ARED advanced resistive exercise device (FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1, FE-2).

Afterwards, Koichi downloaded the exercise data file to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

Gennady had four job items on his discretionary “as time permits” task list for today:-

* Packing & stowing a BVN air heater fan and two VD (vozduchovodiy, air ducts) in the FGB,
* Conducting the periodic audit/inventory of RS (Russian Segment) medical kits (which total about 35),
* Setting up the St. George Ribbons (lentochek) delivered on Progress 33P on an interior panel so that they are visible during TV conferences [the traditional patriotic “St. George’s Ribbons” event observes the 64th Anniversary of the WW2 victory in August], and
* Another session for Russia’s Environmental Safety Agency (EKON), making observations and taking KPT-3 aerial photography of environmental conditions on earth using the Nikon D2X with the SIGMA 300-800mm telephoto lens.

Weekly Science Update (Expedition Nineteen — Week 6)

3-D SPACE: Ongoing.

AgCam (Agricultural Camera): Planned.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): “Sample 5 is now in perfect focus and we’re replacing the ceiling tiles that used to be above Peter Lu (Harvard PI) when he jumped for joy. Thanks for the hard work you’ve put in on samples 8, 9, and 10, the crystal samples. We are reviewing these photos now and will have feedback to give you soon. We have also compiled responses to your requests for feedback with example photos, which you will be receiving soon. Thank you again. We are very grateful for all you’re doing.”

BIOLAB (ESA): “On 5/12, a leak-tightness test of the Rotor B Life Support System (LSS) loop has been successfully carried out, with the support of Koichi to open/close BIOLAB valves. The LSS circuit is confirmed to be leak tight on the BIOLAB facility side. At position B6, a leak has been localised, and ground teams are investigating the next troubleshooting steps after this outcome. On 5/13, ground commanding has been successfully performed to test the Rotor B bellow mechanism. Thanks a lot for your support: this week activities conclude an important part of the preparatory activities for the next BIOLAB experiment.”

Biological Rhythms (JAXA): Ongoing. Two runs are for Barratt, two are planned for De Winne.

BISE (CSA, Bodies in the Space Environment): Ongoing.

BISPHOSPHONATES: “Koichi, thanks for your continued support.”

CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Ongoing.

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack): “The final preparations for starting the FLEX test points were completed this week as Koichi completed a complicated QD troubleshooting procedure and replaced a depleted fuel reservoir inside the CIR combustion chamber. We certainly appreciate the effort it took to confirm the size of the QD leak and verify it was constrained to a specific case when one end of the hose is disconnected. With the replacement of the fuel reservoir, the CIR is ready to fill the chamber with the desired pressure and mixture of gases for the first three FLEX test points. Thanks very much for getting us ready to start our science operations, Koichi!”

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Cosmo Flower mission was completed by Wakata (slipped from I-18)

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the Protein Crystallization Diagnostic Facility (PCDF) experiment. Over the last couple of weeks, some power trips and reboots of EDR and/or PCDF have been encountered, and engineering teams are pin-pointing a flimsy EDR ESEM Board as the root cause of these problems. ESA plans to perform the swap of this ESEM Board with a spare available on orbit in the near-future. To perform the swap, EDR and PCDF will have to be powered down for 1h30m minimum (impact to science to be minimized with respect to planned runs).

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): Operating.

EPM (European Physiology Module): Underway.

EPO (Educational Payload Operations): Reserve.

EPO J-Astro Report (JAXA): “Website has been acquiring a good reputation. We would like to thank you for your hard work and we look forward to seeing more interesting reports from you.”

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): Complete.

EPO Moon Score (JAXA): Planned.

EPO Try Zero-G (JAXA): “Second session of the Try Zero G was successfully finished. Effective arrangement and make the best of inventiveness by crewmembers were appreciated by not just the EPO team but also media people who have visited TKSC for coverage. We also got good HDTV downlink data. JAXA team greatly appreciates the crew and POIC support.”

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): Complete.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): Platform nominal, with the exception of a MIL-Bus error on 5/13. EuTEF platform and its instruments have been safe-configured for the NH3 venting planned on 5/15. – DEBIE-2: On-going science acquisition, also measured (debris) signals during yesterday’s NH3 venting event; – DOSTEL: On-going science acquisition; – EXPOSE: On-going science acquisition. On 5/13, the valves and the lids of the Experimental Trays have been closed for the upcoming NH3 venting event. It is planned to keep those closed for several days (~1-2 weeks), and presently no exact date has been set to re-open them yet; – EVC: Instrument is too cold to be operated for the time being; – FIPEX: All measurement sensors have been switched off for the NH3 venting event. The next EOP is planned to start on 5/17; – MEDET: On 5/6 the instrument was switched off because the instrument was getting too cold. Expected to be turned on when Beta angle improves, possibly before the NH3 venting event. – PLEGPAY: Inactive. Ground teams are preparing for the restart of the Langmuir Probe measurement campaign by Week#7, if IMMT gives a go-ahead.

FACET (JAXA): “Koichi, we would like to thank you for successfully rotating the cassette. After your rotation, optical adjustment was finished by Ground Operation. Now, we can move into second half series of this very successful experiment from next week.”

FSL (Fluid Science Laboratory): On 5/12, FSL was activated to acquire MMA measurements in conjunction with 33P docking. FSL MMA measurements will continue to be performed in the coming weeks to support the PCDF science runs.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter (JAXA): One of two sessions for FE-2 is slipped from Increment 18. Scheduled on 5/21.

HQPC (JAXA): To be launched by 34P.


IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.

Integrated Immune: Complete.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

MAXI (JAXA): Exposed Payload, to be launched on 2JA.

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): Since 5/5 several nucleation runs have been successfully performed with EP1 reactor. With these runs, the science team expects to better understand the influence of temperature on the onset of nucleation process at various super-saturation (i.e. protein solubility) levels. For the EP4 reactor which contains some nice crystals to be returned by 2J/A, science team has decided to progressively decrease its temperature from +23degC to +20degC by 1degC steps (over 24hrs) with some days in-between the temperature decreases.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Thanks for completing the download and initialization of the Sleep Actiwatches on Friday. Two Actiwatches were downloaded, two batteries changed, and four Actiwatches were initialized. The next download and initialization should occur around the second week of June.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): Awaiting the start of the next Sun visibility window which is expected to begin on 5/23. The platform has been put in safe mode for the 33P docking event on 5/12. During yesterday’s (5/15) NH3 venting event, the platform was in idle mode (i.e. motor-controlled and fixed) and all the instrument covers were closed.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

SPICE (Smoke Point In Co-flow Experiment): “Mike and Koichi, We are looking forward to the VolSci activities that are scheduled today (5/16).”

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): “Through 5/12, the ground has received a total of 6,850 of ISS CEO imagery for review and cataloguing. Review of your imagery has been slowed as we support STS-125 imagery review. However we are pleased to confirm your acquisition of views of the following targets: Three Gorges Dam, Yangtze River, China (excellent mapping views of this rarely-photographed target); Key Largo, Florida; Vredefort Impact Crater, South Africa; St. Helena Island, South Atlantic (best-ever Astronaut photography to date); Buenos Aires, Argentina; Montevideo, Uruguay; and Port Desire, Argentina. Way to go crew! Although none of your imagery is being published this week, a number of candidates for future publication have been identified. Thanks for continued support of our payload.”

No CEO (Crew Earth Observation) targets uplinked for today. (In recent days, ISS daylight-awake orbit tracks have shifted rapidly into the Southern Hemisphere which is now some seven weeks into the fall season, and both day length and sun elevation are significantly lowering. This situation along with deteriorating seasonal weather greatly limits good view opportunities for targets. The situation is further aggravated by the fact that the ISS orbit tracks nearly parallel with the terminator. The consequence is very low light right of track, low light near nadir, and adequate to good light left of track. Beginning a few days ago and for the near future, there may be no targets with suitable illumination or weather.)

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 7:47am EDT [= epoch])
Mean altitude — 350.8 km
Apogee height — 357.1 km
Perigee height — 344.5 km
Period — 91.55 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0009412
Solar Beta Angle — 56.6 deg (magnitude peaking)
Orbits per 24-hr. day — 15.73
Mean altitude loss in the last 24 hours — 78 m
Revolutions since FGB/Zarya launch (Nov. 98) — 60096

Significant Events Ahead (all dates Eastern Time, some changes possible!):
05/18/09 — Progress M-01M/32P deorbit (~3:00pm EDT)
05/27/09 — Soyuz TMA-15/19S launch
05/29/09 — Soyuz TMA-15/19S docking (FGB nadir)
Six-person crew on ISS
06/05/09 — Russian EVA-22
06/10/09 — Russian EVA-23
06/13/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
07/17/09 — Progress M-02M/33P undock & deorbit
07/20/09 — Soyuz TMA-14/18S relocation (from SM aft to DC1)
07/24/09 — Progress 34P launch
07/26/09 — Progress 34P docking (SM aft)
08/06/09 — STS-128/Discovery/17A — MPLM (P), LMC
09/01/09 — H-IIB (JAXA HTV-1) launch — tentative
09/07/09 — H-IIB (JAXA HTV-1) berth
09/30/09 — Soyuz TMA-16/20S launch
10/02/09 — Soyuz TMA-16/20S docking (SM aft, until MRM2 w/new port)
10/08/09 — H-IIB (JAXA HTV-1) unberth
10/11/09 — Soyuz TMA-14/18S undock
10/15/09 — Progress 35P launch
11/10/09 — 5R/MRM2 (Russian Mini Research Module) on Proton — tentative
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/07/09 — Soyuz TMA-17/21S launch
12/26/09 — Progress 36P launch
02/03/10 — Progress 37P launch
02/??/10 — STS-130/Endeavour/20A — Node-3 + Cupola — tentative
02/11/10 — STS-131/Atlantis/19A — MPLM(P), LMC — tentative
04/02/10 — Soyuz TMA-18/22S launch
04/08/10 — STS-132/Discovery/ULF4 — ICC-VLD, MRM1 — tentative
04/27/10 — Progress 38P launch
05/29/10 — Soyuz TMA-19/23S launch
05/31/10 — STS-133/Endeavour/ULF5 — ELC3, ELC4 — tentative
06/??/10 — ATV2 — Ariane 5 (ESA)
06/25/10 — Progress 39P launch
08/11/10 — Progress 40P launch
09/29/10 — Soyuz TMA-20/24S launch
12/??/11 — Proton 3R/MLM w/ERA.
10/19/10 — Progress 41P launch
12/??/11 — 3R Multipurpose Laboratory Module (MLM) w/ERA — on Proton

SpaceRef staff editor.